# FLARE: Fusing Language Models and Collaborative Architectures for Recommender Enhancement

### v1

@@ 3.4 Critiquing
We propose a novel masking and evaluation paradigm utilizing critiquing to assess whether hybrid recommenders can appropriately leverage both modalities. This consists of adding a critique in the form of a textual product category $S$ to each item $i$ in the user’s sequence $S$, providing a broad hint to direct the model to the next recommended item. That is, each item $i$ in the user sequence $S$ becomes $i=(ID,T,S)$, where $S$ is the critique string. For masked locations during training and inference, we only mask the $ID$ and item description $T$, leaving $S$ unmasked, which the model can utilize as a hint towards the target item (i.e., $i=([\langle mask\rangle,\langle mask\rangle,S)$).
The key to this task is that hints are textual, therefore requiring recommenders to adapt to a text modality by relating the critique to the other signals in the user’s sequence. Note that this task formulation has applications to many real-life uses of recommender systems, such as refining recommendations based on the user’s natural language query text or the webpage subsection the user is currently navigating. In essence, critiquing strings can be used as a vehicle to steer a model predictions to incorporate the user’s current context.
@@
$T^{\prime}_{k}$ $=LLM(T_{k})+\theta_{T}$ (3)
$S^{\prime}_{k}$ $=LLM(S_{k})+\theta_{S}$ (4)
$t_{k}$ $=Linear(Perceiver([S^{{}^{\prime}}_{k};T^{{}^{\prime}}_{k}]))$ (5)
$c_{k}$ $=t_{k}+e_{k}$ (6)
where $\theta_{T}$ and $\theta_{C}$ are type embeddings to denote text and critiquing as distinct types, “;” denotes concatenation, and “$+$” denotes vector addition, and $e_{k}$ is the embedding for item $ID_{k}$. When masking items for training or during inference, we replace $T^{\prime}_{k}$ and $e_{k}$ at the desired positions with $\langle mask\rangle$ embeddings, allowing the model to utilize $S^{\prime}$ as a hint. By treating critique text as a separate input modality, we hypothesize that the model can learn the connection between critique text and item context. Note that critiquing is applied with Flare through an optional additional input to the Perceiver model, which can be omitted.
3.5 Training
We train our model using MLM loss by masking items in the user’s sequence with 15% probability, ensuring that at least one item is masked. For the total loss, we use a weighted sum of MLM loss and the ID-Text contrastive loss:
$\mathcal{L}_{total}=\alpha\mathcal{L}_{MLM}+(1-\alpha)\mathcal{L}_{c}$ (7)
controlled by the hyper-parameter $\alpha$, which we found setting $\alpha=0.5$ yielded optimal results. For the contrastive loss, we utilize a softmax temperature of $\tau=0.3$ and softmax additive margin term of $m=0.2$.
4 Results
@@
  • RQ2: Can simple models like Flare scale to large item count datasets?
@@
4.1 Datasets
Dataset # Users # Items Avg. Length
Games 55 132 17 389 8.99
Office 101 357 27 932 7.87
Scientific 11 011 5 327 6.97
Music 27 525 10 611 8.4
Arts 56 110 22 855 8.77
Pets 47 569 37 970 8.84
Clothing 1 219 592 376 843 9.2
Table 1: Amazon product reviews dataset statistics
We use datasets from Amazon Product Reviews (ni-etal) to evaluate our model. Table 1 shows the statistics [^1] We observed that the statistics from the downloaded Recformer datasets differ from the statistics reported in the Recformer paper, most notably the Games dataset. for the datasets.
Amazon Product Reviews
@@ 4.1 Datasets
The Amazon Product Reviews dataset contains 233.1 million reviews from May 1996 to October 2018, covering 29 categories ranging from Office Supplies to Video Games. Each review includes item features (title, description, category, price) and review features (text, timestamp, score, summary). User sequences can be created by sorting each user review by timestamp, creating a sequential recommendation task with the goal of predicting what the user will review next.
We compare our results to Recformer (recformer). To enable fair comparison, we use the six dataset splits provided in Recformer: “Video Games”, “Office Supplies”, “Industrial and Scientific”, “Arts, Crafts and Sewing”, “Pet Supplies” and “Musical Instruments.” Each item has at least five reviews, and every user has given at least five reviews. We use the same preprocessing procedure as Recformer, which includes trimming the user sequences to a max length of 51; training on all sequences, regardless of length; and, evaluating on only sequences with 4 or more items for validation and test. Recformer associates and trains each item with the category, title, and brand features.
@@
In addition to the six “small” datasets above, we also evaluate on one medium-sized dataset, “Clothing, Shoes and Jewelry,” which contains 378k items. The preprocessing procedure is the same as before, but rather than trim, we filter all user sequences with more than 50 items and items with a missing title. We associate and train each item to the category and item features. Notably, this dataset is approximately 9x larger than the largest small dataset. Statistics for this dataset too can be found in Table 1.
4.2 Experimental Setup
Baselines
@@ 4.2 Experimental Setup
We compare Flare Text + ID against two classes of state-of-the-art sequential recommenders:
  1. 1.
@@
    ID only models, like Bert4Rec (bert4rec), which utilizes a transformer-based approach with BERT-style (bert) masked language-model training for item sequences. We present updated baseline results after extensive tuning, which we call Flare ID. We emphasize that Flare ID is a Bert4Rec model, with our hyper-parameter tuning.
  2. 2.
    Text-only models, specifically Recformer (recformer), which departs from using explicit item ID representations and proposes a text-based representation using learned key-value pairs.
  3. 3.
@@
    Hybrid Text + ID models, specifically S3Rec (s3rec), which utilizes self-supervised pre-training tasks to align items and attributes through mutual information maximization.
Evaluation Settings
Params Games Office Scient. Music Arts Pets
# Layers 2 2 4 8 2 2
# Heads 2 2 16 8 8 16
Model Dims 64 64 512 256 768 1024
Hidden Dims 256 256 2048 1024 3072 4096
Learning Rate 1e-3 1e-3 1e-4 1e-5 1e-4 1e-4
Batch 1 1 16 16 16 32
Total Steps
50K 50K 50K 50K 50K 10K
Table 2: Hyper-Parameters for Flare ID (tuned Bert4Rec)
Params Games Office Scient. Music Arts Pets
# Layers 8 2 2 2 4 2
# Heads 16 4 8 2 4 8
Model Dims 1024 768 256 1024 512 256
Hidden Dims 4096 3072 1024 4096 2048 1024
Learning Rate 1e-4 1e-4 1e-4 1e-5 1e-4 1e-4
Batch 2 16 8 2 8 16
Total Steps
5K 5K 5 25K 10K 25K
Perceiver #Heads
16 16 8 8 2 8
Perceiver #Layers
8 2 2 8 2 2
Perceiver #Latents
2 8 2 4 2 2
Weight Decay
1e-3 1e-2 1e-3 1e-3 1e-3 1e-3
Table 3: Hyper-Parameters for for Flare T+ID
@@
Following prior work (recformer; lc-rec; bert4rec), we adopt the leave-one-out strategy for evaluation. We reserve each user’s penultimate and last items for validation and test, respectively, using the rest of the user sequence for train.
@@
Since training sequences can be of varying length (Table 1), we pack multiple training sequences into a single input example such that each packed example is as close to the same size as possible. Then, during training, we restrict attention to just the items within the same sequence. This allows us to maximize the throughput of our hardware and avoid wasteful computation due to excessive padding tokens. As a result, the effective batch size during training can vary from step to step, depending on the length of the sequences.

### v2

@@ 3.4. Critiquing
We propose a novel masking and evaluation paradigm using critiquing to assess whether hybrid recommenders can appropriately leverage both modalities. This consists of adding a critique in the form of a textual item category $S$ to each item $i$ in the user’s sequence $U$, providing a broad hint to direct the model to the next recommended item. That is, each item $i$ in the user sequence $U$ becomes $i=(ID,T,S)$, where $S$ is the critique string. For masked locations during training and inference, we only mask the $ID$ and item text $T$, leaving $S$ unmasked, which the model can utilize as a hint towards the target item (i.e., $i=([\langle mask\rangle,\langle mask\rangle,S)$).
The key to this task is that hints are textual, therefore requiring recommenders to adapt to a text modality by relating the critique to the other signals in the user’s sequence. Note that this task formulation has applications to many practical uses of recommender systems, such as refining recommendations based on the user’s natural language query text or the webpage subsection the user is currently navigating. In essence, critiquing strings can help steer a model’s predictions to incorporate the user’s current context.
@@
(3) $T^{\prime}_{k}$ $=LLM(T_{k})+\theta_{T}$
(4) $S^{\prime}_{k}$ $=LLM(S_{k})+\theta_{S}$
(5) $t_{k}$ $=Linear(Perceiver([S^{{}^{\prime}}_{k};T^{{}^{\prime}}_{k}]))$
(6) $c_{k}$ $=t_{k}+e_{k}$
where $\theta_{T}$ and $\theta_{C}$ are type embeddings to denote text and critiquing as distinct types, “;” denotes concatenation, “$+$” denotes vector addition, and $e_{k}$ is the embedding for item $ID_{k}$. When masking items for training or during inference, we replace $T^{\prime}_{k}$ and $e_{k}$ at the desired positions with $\langle mask\rangle$ embeddings, allowing the model to utilize $S^{\prime}$ as a hint. By treating critique text as a separate input modality, we hypothesize that the model can learn the connection between critique text and item context. Note that critiquing is applied with Flare through an optional additional input to the Perceiver model, which can be omitted.
3.5. Training
We train Flare using the MLM loss by masking items in the user’s sequence with 15% probability, ensuring that at least one item is masked. For the total loss, we use a weighted sum of MLM loss and the ID-Text contrastive loss:
(7) $\mathcal{L}_{total}=\alpha\mathcal{L}_{MLM}+(1-\alpha)\mathcal{L}_{c}$
controlled by the hyper-parameter $\alpha$, which we found $\alpha=0.5$ yielded optimal results. For the contrastive loss, we use a softmax temperature $\tau=0.3$ and softmax additive margin term $m=0.2$.
4. Experimental Results
@@
  • RQ2: Can hybrid models like Flare perform well on larger item count datasets?
@@
4.1. Datasets
Table 1. Statistics for the Recformer- and CALRec-processed Amazon Product Reviews dataset. I/U is the average number of items per user, or average sequence length. Note that the Clothing dataset preprocessing follows that of Recformer, except that it filters out long sequences rather than trimming them (Section 4.1.2).
Recformer CALRec
Dataset # Users # Items I/U # Users # Items I/U
Games 55,132 17,389 9.00 50,940 17,383 8.97
Office 101,357 27,932 7.88 44,736 27,482 7.87
Scientific 11,011 5,327 6.98 9,461 5,282 7.04
Music 27,525 10,611 8.40 25,577 10,599 8.39
Arts 56,110 22,855 8.77 47,197 22,828 8.72
Pets 47,569 37,970 8.84 43,135 37,712 8.82
Clothing 1,219,592 376,843 9.20
@@ 4.1. Datasets
We use datasets from Amazon Product Reviews (Ni et al., 2019) to evaluate our model. The Amazon Product Reviews dataset contains 233.1 million reviews from May 1996 to October 2018, covering 29 categories ranging from Office Supplies to Video Games. Each review includes item features (title, description, category, price) and review features (text, timestamp, score, summary). User sequences are created by grouping and sorting reviews by timestamp, enabling a sequential recommendation task to predict the next item the user reviews.
We benchmark against Recformer (Li et al., 2023) and CALRec (Li et al., 2024), which report the current state-of-the-art results. To meaningfully compare our results to each of those projects, we use the same Amazon Product Reviews dataset preprocessing procedure that each of those projects use. Where available, we either downloaded and trained with the relevant preprocessed datasets or ran the project’s pipeline to preprocess the raw dataset. Dataset statistics are in Table 1.
4.1.1. Recformer-Processed Dataset
We use the six dataset splits provided in Recformer: [^1] We observed that the statistics from the downloaded Recformer datasets differ from the statistics reported in the Recformer paper, most notably the Games dataset. “Video Games”, “Office Supplies”, “Industrial and Scientific”, “Arts, Crafts and Sewing”, “Pet Supplies” and “Musical Instruments.”
4.1.2. Clothing Dataset
@@
In addition to the six small datasets above, we also evaluate on one medium-sized dataset, “Clothing, Shoes and Jewelry,” which contains 378k items. The preprocessing procedure is the same as Recformer, but rather than trim, we filter all user sequences with more than 50 items and items with a missing title. Notably, this dataset is approximately 9x larger than the largest small dataset.
4.1.3. CALRec-Processed Dataset
CALRec preprocessing notably performs data-deduplication on user sequences. We follow CALRec’s deduplication methodology and remove consecutive items by the same user that have the same item ID (ASIN number), review text, rating, and timestamp. The motivation for deduplication is to mitigate trivial recommendation strategies like always recommending the most recent item.
4.2. Experimental Setup
4.2.1. Baselines
@@
We compare Flare against three classes of state-of-the-art sequential recommenders.
@@
First are ID only models, like Bert4Rec (Sun et al., 2019), which utilize a transformer-based approach with BERT-style (Devlin et al., 2019) masked language-model training for item sequences. We show updated baselines after extensive tuning, which we call Bert4Rec. We emphasize that Bert4Rec is Bert4Rec, with our hyperparameter tuning.
Second are text-only models, which use a text-based representation of user history and items without an explicit item ID representation. We focus benchmarking against Recformer (Li et al., 2023) and CALRec (Li et al., 2024), which report the current state-of-the-art results.
@@
Finally, we compare against hybrid Text + ID models, specifically (a) S3Rec (Zhou et al., 2020), which utilizes self-supervised pre-training tasks to align items and attributes through mutual information maximization; and (b) UniSRec (Hou et al., 2022), which relies on text-based item representations but can incorporate IDs in a fine-tuning step.
One notable difference between Flare and the prior models Recformer and CALRec is that Flare trains a separate model for each dataset. In contrast, Recformer and CALRec propose a two-stage training approach, in which a single model is pre-trained on multiple datasets, followed by a fine-tuning stage on reported evaluation datasets. We view this as a difference in model training approach.
4.2.2. Evaluation Settings
@@
Following prior work (Li et al., 2023; Zheng et al., 2024; Sun et al., 2019), we adopt the leave-one-out strategy for evaluation. We reserve each user’s penultimate and last items for validation and test, respectively, using the rest of the sequence for training.
@@
Since training sequences can be of varying length (Table 1), we pack multiple training sequences into a single input example such that each packed example is as close to the same size as possible. Then, during training, we restrict attention to just the items within the same sequence. This maximizes the throughput of our hardware and avoids wasteful computation due to excessive padding tokens. As a result, the effective batch size during training can vary from step to step, depending on the length of the sequences.